The net revenue of a company rose 3.6 %, while inflation rate was 3.0 %. Calculate the relative real change in revenue as percentage.

 


1
Expert's answer
2022-02-14T18:04:29-0500

Let revenue of a particular year is "' x '"

then revenue of next year will be

"\begin{aligned}

& x+3.6 \% \text { of } x=x+0.036 x \\

=& 1.036 x

\end{aligned}"

now, Inflation 3%

So,

"\begin{aligned}

&=1.036 x-1.036 x \times 3 \% \\

&=1.036 x-0.03108 x \\

&=1.00492 x

\end{aligned}"

Revenue of particular year was "x" after one year value of revenue is "1.00492 x" .

So, change "\%=\frac{1.00492 x-x}{x} \times 100 \%=0.492 \%"

Relative real change in revenue "=0.492 \%"